OpenStudy (anonymous):

What is the log(base e) of (1/3) e^3a

OpenStudy (anonymous):

\[\log_{e} ((1/3)e ^{3a})\]

OpenStudy (kropot72):

\[\ln (\frac{1}{3}\times e ^{3a})=\ln \frac{1}{3}+3a\]